# LBMamba: Locally Bi-directional Mamba

### v1

@@ Appendix D Implementation details of WSI experiments
Feature extractor. We use UNI (Chen et al., 2024), a well-known and current SOTA foundation model for feature extraction, which is a ViT-L/16 pretrained on more than 100 million pathology patches from from over 100,000 H&E-stained WSIs across 20 major tissue types. UNI is pretrained in a self-supervised manner using DINOv2 (Oquab et al., 2023).
WSI pre-processing. We extract patches from WSIs at 20x magnification with no overlapping. The patch size is set to 512x512 pixels. We used the preprocessing tool in CLAM (Lu et al., 2021) to segment and extract tissue regions.
Training. We use AdamW (Loshchilov & Hutter, 2017) to optimize the models for $20$ epochs of training with batch size being 1. The initial learning rate is $0.0001$ and we use cosine annealing decay to adjust it.
Appendix E Details of WSI datasets
Prostate cancer grading based on PANDA. Panda dataset (Bulten et al., 2022) consists of 10,614 biopsies of prostate cancer. Its labels are their ISUP grading, totally 6 categories: grade 0 (2890 slides), grade 1 (2666 slides),grade 2 (1343 slides), grade 3 (1242 slides), grade 4 (1249 slides), and grade 5 (1224 slides). We random split PANDA into 80:10:10 train/validation/test sets.
Breast invasive carcinoma subtyping on TCGA-BRCA. TCGA-BRCA comprise 1033 H&E WSIs with 2 subtypes: invasive ductal carcinoma (822 slides) and invasive lobular carcinoma (211 slides). We follow (Chen et al., 2022) to get the train/validation/test set with 841:96:96 slides.
Non-small cell lung carcinoma subtyping on TCGA-NSCLC. The dataset include 957 H&E breast carcinoma WSIs, including 2 subtypes: lung adenocarcinoma (490 slides) and lung squamous cell carcinoma (468 slides). We follow (Chen et al., 2022) to split the dataset into train/validation/test set with 785:86:87 slides.

### v2

@@ Appendix E Implementation details of WSI experiments
Feature extractor. We use UNI (chen2024uni), a well-known and current SOTA foundation model for feature extraction, which is a ViT-L/16 pretrained on more than 100 million pathology patches from from over 100,000 H&E-stained WSIs across 20 major tissue types. UNI is pretrained in a self-supervised manner using DINOv2 (oquab2023dinov2).
WSI pre-processing. We extract patches from WSIs at 20x magnification with no overlapping. The patch size is set to 512x512 pixels. We used the preprocessing tool in CLAM (lu2021data_clam) to segment and extract tissue regions.
Training. We use AdamW (loshchilov2017decoupled_adamw) to optimize the models for $20$ epochs of training with batch size being 1. The initial learning rate is $0.0001$ and we use cosine annealing decay to adjust it.
Appendix F Details of WSI datasets
Prostate cancer grading based on PANDA. Panda dataset (bulten2022artificial_panda) consists of 10,614 biopsies of prostate cancer. Its labels are their ISUP grading, totally 6 categories: grade 0 (2890 slides), grade 1 (2666 slides),grade 2 (1343 slides), grade 3 (1242 slides), grade 4 (1249 slides), and grade 5 (1224 slides). We random split PANDA into 80:10:10 train/validation/test sets.
Breast invasive carcinoma subtyping on TCGA-BRCA. TCGA-BRCA comprise 1033 H&E WSIs with 2 subtypes: invasive ductal carcinoma (822 slides) and invasive lobular carcinoma (211 slides). We follow (chen2022scaling_hipt) to get the train/validation/test set with 841:96:96 slides.
Non-small cell lung carcinoma subtyping on TCGA-NSCLC. The dataset include 957 H&E breast carcinoma WSIs, including 2 subtypes: lung adenocarcinoma (490 slides) and lung squamous cell carcinoma (468 slides). We follow (chen2022scaling_hipt) to split the dataset into train/validation/test set with 785:86:87 slides.
Appendix G Qualitative evaluation on segmentation and detection tasks
To complement the quantitative metrics of the segmentation and detection tasks presented in section 4.3, we provide a qualitative evaluation of our models on them. Figure 6 displays a qualitative comparison for semantic segmentation on the ADE20K dataset. The visualizations show that the masks produced by LBVim-Ti model are slightly less precise than those of the Vim-Ti baseline. However, the advantage of our approach becomes evident with the scaled-up LBVim-300 model. It generates visibly superior segmentation masks with cleaner boundaries and more accurate class assignments, corroborating the significant mIoU improvements reported in section 4.3. Similarly, figure 7 provides a visual comparison for object detection. In this task, LBVim-Ti achieves performance comparable to the Vim-Ti baseline, successfully identifying the main objects in the scene. Once again, the scaled-up LBVim-300 model demonstrates a clear leap in performance, producing more accurate detections. This qualitative improvement directly aligns with the higher AP scores reported in section 4.3, confirming the effectiveness of our method for detection tasks.
Figure 6: Qualitative segmentation comparison on the ADE20K dataset. While LBVim-Ti produces slightly less precise masks than the Vim-Ti baseline, the scaled-up LBVim-300 demonstrates the strength of our architecture with visibly superior segmentation quality.
Figure 7: Qualitative comparison of detection results from Vim-Ti, LBVim-Ti, and LBVim-300 on the ADE20K dataset. LBVim-Ti achieves performance comparable to the Vim-Ti baseline, while our scaled-up LBVim-300 model produces visibly more accurate detection.